5(x-3)+4(2x+2)=4(x+2)+3

Show how you got the answer

Answer by LVSMath(2)

Distribute:5x-15+8x+8=4x+8+3
Combine like terms: 13x-7=4x+11
Simplify: 13x-7+7=4x+11+7
13x=4x+18
13x-4x=4x-4x+18
9x=18
Divide: 9x/9=18/9
x=2